OpenStudy (anonymous):

Decide whether the proportion is true or false by writing the ratios in the lowest terms. 65/25=78/12

OpenStudy (anonymous):

65/25=13/5......divided by 5 and 78/12=13/2....... divided by 6 false proportion

OpenStudy (anonymous):

Thank you